# PowerFlowMultiNet: Multigraph Neural Networks for Unbalanced Three-Phase Distribution Systems

### v2

@@ Results
We summarize the performance of our approach in Table III. PowerFlowMultiNet is always more precise than FCNN, and even 100 times more precise on the larger 123-bus topology. Overall, our approach achieves very low error compared to the OpenDSS ground truth.
Next, we evaluate the efficiency of our approach in Figure 3. Our approach is 100 times faster than Model-based, with a mean execution time of $67.75e-$ 6 on 123-bus compared to $46.52e-$ 4 required by model-based. More notably, our approach achieves stable execution time between 13-bus and 123-bus cases and a very low standard deviation, confirming its stability and scalability compared to model-based. Our approach behaves similarly even for cases hard to optimize by model-based.
Comparison of relative error of PowerFlowMultiNet and FCNN with load perturbation of ±$\pm 10\%$ and ±$\pm 50\%$.
Load Approach Case $P_{MEAN(STD)}$ $Q_{MEAN(STD)}$
±$\pm 10\%$ OURS 13-Bus $5.46e^{-5}$($9.57e^{-5}$) $2.80e^{-4}$($4.48e^{-4}$)
123-Bus $5.77e^{-7}$($1.00e^{-6}$) $7.29e^{-7}$($1.11e^{-6}$)
FCNN 13-Bus $7.96e^{-4}$($1.44e^{-3}$) $8.57e^{-4}$($1.28e^{-3}$)
123-Bus $1.24e^{-4}$($1.89e^{-4}$) $1.25e^{-4}$($1.82e^{-4}$)
±$\pm 50\%$ OURS 13-Bus $1.21e^{-3}$($2.45e^{-3}$) $5.75e^{-3}$($1.08e^{-2}$)
123-Bus $9.13e^{-6}$($1.69e^{-5}$) $1.33e^{-5}$($2.10e^{-5}$)
FCNN 13-Bus $4.15e^{-3}$($1.38e^{-2}$) $8.09e^{-3}$($1.61e^{-2}$)
123-Bus $6.97e^{-4}$($1.12e^{-3}$) $7.75e^{-4}$($1.28e^{-3}$)
Figure 3: CPU Run time for our approach and OpenDSS
Conclusion
This paper is the first to tackle the unbalanced three-phase power flow problem as a multi-graph neural network. Our framework, PowerFlowMultiNet, includes a novel graph embedding of unbalanced distribution grids and an efficient GNN training algorithm for power flow prediction. Our preliminary empirical study on 13-bus and 123-bus distribution grids demonstrates both its effectiveness compared to traditional Deep learning methods and its efficiency compared to Model-based methods. We believe our work paves the way for novel applications of GNN for distribution grid optimizations and operations.

### v3

@@ Results
We summarize the performance of our approach in Table I. The gray cells indicate the lowest errors per scenario. PowerFlowMultiNet is always more precise than the FCNN and GNN models (except for voltage magnitude with small grids and 10% load perturbation, where GNN achieves as low errors as our approach), while having execution costs similar to simpler approaches. Overall, our approach achieves very low error compared to the OpenDSS ground truth.
Next, we evaluate the efficiency of our approach. Our approach is up to 15 times faster than Model-based on a 13-bus grid (e.g., for 10% load variation), up to 37 times faster on 123-bus grid (e.g., for timeseries load variation), and up to 38 times faster on 906-bus (e.g., for 50% load variation).
Notably, our approach achieves stable execution time for all grid and a very low standard deviation, confirming its stability and scalability to larger grids. Our approach behaves similarly even for cases hard to optimize by model-based approaches.
IV Conclusion
This paper is the first to tackle the unbalanced three-phase power flow problem as a multigraph neural network. Our framework, PowerFlowMultiNet, includes a novel graph embedding of unbalanced distribution grids and an efficient GNN training algorithm for power flow prediction. Our empirical study on 13, 123, and 906 bus distribution grids demonstrates its effectiveness compared to existing deep learning methods and its efficiency compared to model-based methods. We believe that our work paves the way for novel applications of GNN for distribution grid optimizations and operations.